OpenStudy (clamin):

Which of the following pairs of functions are inverses of each other? a. f(x)= x/6 + 8 and g(x)= 6x-8 b. f(x)= 5x-11 and g(x) = x+11/ 5 c. f(x)= 2+ 3 square root x and g(x)= 2-x^3 d. f(x)= 7/x - 9 and g(x) =x+9/7

OpenStudy (clamin):

@Jaynator495 @pooja195 @Marie1738 @mathstudent55 @bunnielover948

OpenStudy (fortytherapper):

Do you know how to find inverses? We can go over one if you want. Any answer choice you wanna try out?

OpenStudy (solomonzelman):

In general, \(\color{#000000}{ \displaystyle f^{-1}(x)=g(x) }\) iff \(\color{#000000}{ \displaystyle f\left[g(x)\right]=g\left[f(x)\right]= x }\).

OpenStudy (clamin):

My answer is B @fortytherapper

OpenStudy (clamin):

@ganeshie8